OpenStudy (anonymous):

adding, subtracting, multiplying radicals -3 squareroot 12+3 squareroot 3+3 squareroot 20

OpenStudy (anonymous):

\[-3\sqrt{12}+3\sqrt{3}+3\sqrt{20}\]

OpenStudy (anonymous):

\[-2\sqrt{45}-3\sqrt{20}-2\sqrt{6}\]

OpenStudy (anonymous):

how did u get that answer

OpenStudy (anonymous):

thnks